# Training Language Models on Synthetic Edit Sequences Improves Code Synthesis

### v2

@@ B.1 Prompting
The primary goal of this paper is to introduce a method for re-factorizing code synthesis with LMs by finetuning them on synthetic instruction data. As a result, we evaluate all models using minimal prompt formats, performing no prompt tuning (see Figures 9 and 10). Examples of the prompt formats that we use during evaluation are shown in Figure 8.
@@
To evaluate models on HumanEval, we use both the default “Python version” prompt format in the original benchmark dataset, where a natural language program description is provided to an LM within a docstring, as well as the equivalent, fully natural language prompt format from HumanEvalPack (Muennighoff et al., 2023). The latter format is similar to the structure of the instructions in our finetuning datasets.
To evaluate models on MBPP(+), we use the default prompts from the MBPP benchmark dataset, formatted with specification of the target function name and arguments both inside and outside of the natural language instruction, as shown in Figure 8.
During LM benchmark evaluations, we test models on each of the format variants described above and report scores on the better performing variant for each benchmark only.
B.2 Generation and Parsing
During generation, we continue decoding until an end-of-sequence token is output by an LM. We treat all LM outputs as either Python code or sequences of Python code edits, depending on whether an LM was finetuned on standard instruct or LintSeq instruct data. In the latter case, we post-process outputs by resolving the output edit sequences using the procedure described in Appendix A.2.
B.3 Evaluating Model Checkpoints
B.3.1 Philosophy
@@
B.3.2 Computing Coverage (Pass@K)
Our goal is to probe whether re-parameterizing code synthesis with edit sequences can improve the expressivity of smaller LLM representations, boosting benchmark coverage as a function of samples-per-problem. Hence, we primarily compare finetuned models by evaluating them with the procedures described above on HumanEval and MBPP(+) at a high temperature and a large sample count, computing pass@k for k$\ \in\{1,5,10,20,50\}$ with $N=50$ samples [^2] These are the largest sample counts that are feasible to compute on our hardware given the scope of our experiments. at temperature 1, top-p 0.95. We compute pass@k statistics with the same procedure as Chen et al. (2021). The results of these evaluations are reported throughout the paper and shown in Figures 4, 5 and Tables 8, 9, 10, 11. In each of these Figures and Tables, we identify the most performant checkpoint from each model-data finetuning run by comparing pass@50 score at temperature 1 on HumanEval and MBPP(+) across checkpoints.
Many existing state-of-the-art code synthesis LMs only report temperature-tuned pass@k scores on HumanEval, including Codex, AlphaCode, and Codegen-Mono (Chen et al., 2021; Li et al., 2022b; Nijkamp et al., 2022). Thus, in Tables 1 and 2, we select the single best overall checkpoint of each TinyCodeLM model by averaging HumanEval and MBPP(+) pass@50 score. Then, we temperature-tune each checkpoint’s pass@1 and pass@10 scores when reporting results. On HumanEval, we test temperatures $\tau\in\{0.0,0.2,0.4,0.8,1.0\}$. On MBPP(+), we sweep over a smaller temperature range, $\tau\in\{0.0,0.1,1.0\}$. We perform the same temperature tuning procedure when reporting external model benchmark scores as well, i.e. the scores annotated with “()($\dagger$)( † )” in Tables 1 and 2. When running benchmark evaluations with these external code LMs, we stray from the prompt formatting, generation, and parsing procedures described in Appendices B.1 and B.2; instead, in the interest of a fair evaluation, we reproduce the conventions reported by model authors to report other scores.
Appendix C Pretraining
@@
C.1 Model Architectures and Pretraining Hyperparameters
@@ C.2 Pretraining Data Mix
Appendix D Instruction Finetuning
D.1 Baseline Instruction Dataset
Table 5 displays the data sources that are used to prepare the dataset described in Section 3.2. These data are pooled and preprocessed into instruction-program pairs by stripping away Markdown formatting and natural language explanations from completions (Figure 9 and 10). In our experiments, we use the resultant data to finetune baseline models, comparing their performance to those of LMs finetuned on edit sequences generated with LintSeq from the same set of instruction-program pairs.
@@ D.1 Baseline Instruction Dataset
D.2 Procedures and Hyperparameters
We instruction finetune all models with Microsoft DeepSpeed using the ZeRO++ protocol for stage three sharding. For the largest of these models, we also use CPU parameter offloading to accelerate experiments (Wang et al., 2023a; Ren et al., 2021). When finetuning models on LintSeq data, we add a new token “<|diff|>” to tokenizers (Section 2.4) and resize model embeddings accordingly.
@@ D.2 Procedures and Hyperparameters
Across all of the finetuning experiments conducted in this paper, we train model-data variants with the same batch size and for an equal number of total optimizer steps. This optimizer step count corresponds to ten epochs of finetuning with the baseline instruction tuning dataset described in Section 3.2. We save intermediate checkpoints at equal optimizer step intervals in all experiments.
@@
Table 7: All other instruction finetuning settings, re-used across experiments.
Appendix E More on Synthetic Data Generation with LintSeq
E.1 Examples of Generated Synthetic Edit Trajectories
@@
E.2 Tuning LintSeq Example Count
@@
Appendix F Additional Results
F.1 Pretraining TinyCodeLM
Figure 12: Evaluating the zero-shot Python synthesis capabilities of TinyCodeLM-150M during pretraining on HumanEval and MBPP(+).
Figure 13: Evaluating the zero-shot Python synthesis capabilities of TinyCodeLM-400M during pretraining on HumanEval and MBPP(+).
F.2 Finetuning TinyCodeLM
F.2.1 Inference-Time Scaling Laws
Table 8: HumanEval fixed-temperature coverage scaling results achieved by finetuning TinyCodeLM models (zero-shot, temperature = 1, top-p = 0.95).
HumanEval
tinycodeLM-Instruct 150M - 6.2 10.4 12.3 14.5 18.3
tinycodeLM-RandSeqInstruct 150M 4.0 9.7 11.9 14.4 17.7 (-0.6)
tinycodeLM-LintSeqInstruct 150M 6.4 14.3 17.7 20.4 22.6 (+4.3)
tinycodeLM-Instruct 400M - 6.8 11.4 13.7 16.2 18.9
tinycodeLM-RandSeqInstruct 400M 7.2 12.7 15.6 18.6 22.0 (+3.1)
tinycodeLM-LintSeqInstruct 400M 7.2 14.9 18.2 21.5 26.8 ( +7.9)
Table 9: MBPP(+) fixed-temperature coverage scaling results achieved by finetuning TinyCodeLM models (zero-shot, temperature = 1, top-p = 0.95).
MBPP(+)
tinycodeLM-Instruct 150M - 7.3 16.4 20.4 24.6 30.6
tinycodeLM-RandSeqInstruct 150M 4.3 13.1 18.4 23.7 30.2 (-0.4)
tinycodeLM-LintSeqInstruct 150M 7.5 18.2 23.4 28.5 34.5 (+3.9)
tinycodeLM-Instruct 400M - 9.3 17.9 22.2 26.8 32.4
tinycodeLM-RandSeqInstruct 400M 5.5 15.9 21.5 27.0 34.5 (+2.1)
tinycodeLM-LintSeqInstruct 400M 13.1 25.2 29.9 34.4 39.6 (+7.2)
F.3 Finetuning Gemma 2, Phi-3, and Llama 3.1
F.3.1 Inference-Time Scaling Laws
Table 10: HumanEval fixed-temperature coverage scaling results achieved by finetuning Gemma 2, Phi-3, and Llama 3.1 models (zero-shot, temperature = 1, top-p = 0.95).
HumanEval
Model Variant Parameters pass@1 pass@5 pass@10 pass@20 pass@50
Gemma-2-Instruct 2.6B 16.1 26.3 30.5 34.6 40.2
Gemma-2-LintSeqInstruct 2.6B 21.6 35.0 41.2 48.3 58.5 (+18.3)
Phi-3-Mini-Instruct 3.8B 40.1 48.8 52.1 55.0 57.3
Phi-3-Mini-LintSeqInstruct 3.8B 38.3 63.6 72.0 79.0 86.6 (+29.3)
Llama-3.1-Instruct 8B 34.2 50.3 55.6 59.7 63.4
Llama-3.1-LintSeqInstruct 8B 37.6 61.1 69.2 75.3 82.3 (+18.9)
Phi-3-Med-Instruct 14B 52.8 67.9 72.0 75.0 77.4
Phi-3-Med-LintSeqInstruct 14B 44.7 73.5 80.2 84.5 90.2 (+12.8)
Table 11: MBPP(+) fixed-temperature coverage scaling results achieved by finetuning Gemma 2, Phi-3, and Llama 3.1 models (zero-shot, temperature = 1, top-p = 0.95).
MBPP(+)
Model Variant Parameters pass@1 pass@5 pass@10 pass@20 pass@50
Gemma-2-Instruct 2.6B 22.8 34.2 37.8 41.2 44.6
Gemma-2-LintSeqInstruct 2.6B 27.7 41.3 46.3 50.5 54.7 (+10.1)
Phi-3-Mini-Instruct 3.8B 35.2 45.9 49.1 52.0 56.1
Phi-3-Mini-LintSeqInstruct 3.8B 39.3 57.5 62.4 65.9 69.4 (+13.3)
Llama-3.1-Instruct 8B 38.4 50.2 53.8 56.6 59.4
Llama-3.1-LintSeqInstruct 8B 38.5 56.5 61.6 65.7 69.8 (+10.4)
Phi-3-Med-Instruct 14B 40.3 51.5 54.2 56.3 59.0
Phi-3-Med-LintSeqInstruct 14B 40.0 60.9 67.0 71.5 75.9 (+16.9)
F.4 Computing HumanEval Coverage vs Cumulative Inference-Time FLOPs
In Figure 1, we plot HumanEval coverage as a function of cumulative inference-time FLOPs, comparing the performance and total cost of repeatedly sampling from our instruction finetuned Phi-3 and Llama 3.1 models vs sampling a single generation per problem from very large models like Llama 3.1 405B (Dubey et al., 2024) and Nemotron 4 340B (Adler et al., 2024).
We use the approximations below, drawn from Kaplan et al. (2020), to conservatively estimate the cumulative inference costs of synthesizing solutions to all of the $164$ HumanEval benchmark problems across different models. The models that we compare are all dense transformers, where the majority of the parameters are used in matrix multiplications.
$\approx\text{FLOPs per token}\cdot T_{\text{avg-total-tokens-per-sample}}\cdot K_{\text{samples}}\cdot 164$
For our instruction finetuned models, we determine the quantities $T_{\text{avg-total-tokens-per-sample}}$ by computing token counts over all sets of samples per problem that we obtained to compute the coverage statistics in Figure 4 and Table 10 above. These token statistics are provided in the table below.
Table 12: HumanEval total tokens per sample for finetuned Gemma 2, Phi-3, and Llama 3.1 models (zero-shot, temperature $=1$, top-p $=0.95$). These counts reflect prompt and completion tokens. They are computed using the same samples whose coverage statistics are reported in Table 10.
Model Variant Parameters Avg. Total Tokens Per HumanEval Sample ($n=50$)
Gemma-2-Instruct 2.6B 172
Gemma-2-LintSeqInstruct 2.6B 205
Phi-3-Mini-Instruct 3.8B 218
Phi-3-Mini-LintSeqInstruct 3.8B 247
Llama-3.1-Instruct 8B 178
Llama-3.1-LintSeqInstruct 8B 201
Phi-3-Med-Instruct 14B 230
Phi-3-Med-LintSeqInstruct 14B 247
Note that edit sequence (i.e. LintSeqInstruct finetuned) LMs have slightly higher average token counts per sample due to presence of “diff” descriptor tokens in generations (see Appendix A).
We report zero-shot HumanEval coverage for external models by using the evaluation statistics from Dubey et al. (2024) (Table 18, column one). To estimate cumulative inference-time FLOPs for these models, we employ the approximation expression above and estimate $T_{\text{avg-total-tokens-per-sample}}\approx 200$, reflecting an ensemble over the per sample token counts of standard “Instruct” finetuned models shown in Table 12. Note that this ensembled statistic reflects program generations without chain-of-thought only. As a result, we believe it to be a conservative estimator.

### v3

@@ C.1 Prompting
The primary goal of this paper is to introduce a method for re-factorizing code synthesis with LMs by fine-tuning them on synthetic instruction data. As a result, we evaluate all models using minimal prompt formats, performing no prompt tuning (see Figures 9 and 10). Examples of the prompt formats that we use during evaluation are shown in Figure 8.
@@
To evaluate models on HumanEval, we use both the default “Python version” prompt format in the original benchmark dataset, where a natural language program description is provided to an LM within a docstring, as well as the equivalent, fully natural language prompt format from HumanEvalPack (Muennighoff et al., 2023). The latter format is similar to the structure of the instructions in our fine-tuning datasets. We report results on the prompt format that yields the best score for each model.
To evaluate models on MBPP(+), we use the default prompts from the MBPP benchmark dataset, formatted with specification of the target function name and arguments both inside and outside of the natural language instruction, as shown in Figure 8. As on HumanEval, we report results on the prompt format that yields the best score for each model.
To evaluate models on BigCodeBench(Instruct) and CodeContests, we simply prompt models with the problem descriptions introduced in the original version of the benchmark (Zhuo et al., 2024; Li et al., 2022b).
Finally, to evaluate models on DS-1000, we use the completion format, with precisely the same prompt structures as those used by Wei et al. (2024b).
C.2 Generation and Parsing
During generation, we continue decoding until an end-of-sequence token is output by an LM. We treat all LM outputs as either Python code or sequences of Python code edits, depending on whether an LM was fine-tuned on standard instruct or LintSeq instruct data. In the latter case, we post-process outputs by resolving the output edit sequences using the procedure described in Appendix B.2.
C.3 Evaluating Model Checkpoints
C.3.1 Philosophy
@@
C.3.2 Computing Pass@K
Our goal is to probe whether re-parameterizing code synthesis with edit sequences can improve the expressivity of smaller LM representations, boosting benchmark scores as a function of total test-time compute. Hence, we primarily compare fine-tuned models by evaluating them with the procedures described above across multiple pass@k. We compute unbiased pass@k statistics with the same procedure as Chen et al. (2021). The results of these evaluations are reported throughout the paper.
C.4 Comparing TinyCodeLMs to existing Models in Table 1
Many existing state-of-the-art code synthesis LMs only report temperature-tuned pass@k scores on HumanEval, including Codex, AlphaCode, and Codegen-Mono (Chen et al., 2021; Li et al., 2022b; Nijkamp et al., 2022). Thus, in Table 1, we temperature-tune TinyCodeLM models’ pass@1 and pass@10 scores when reporting results. On HumanEval, we test temperatures $\tau\in\{0.0,0.2,0.4,0.8,1.0\}$. On MBPP(+), we sweep over a smaller temperature range, $\tau\in\{0.0,0.1,1.0\}$. We perform the same temperature tuning procedure when reporting external model benchmark scores as well, i.e. the scores annotated with “()($\dagger$)( † )” in Table 1. When running benchmark evaluations with these external code LMs, we stray from the prompt formatting, generation, and parsing procedures described in Appendices C.1 and C.2; instead, in the interest of a fair evaluation, we reproduce the conventions reported by model authors to report other scores.
Appendix D Pretraining
@@
D.1 Model Architectures and Pretraining Hyperparameters
@@ D.2 Pretraining Data Mix
Appendix E Instruction Fine-tuning
E.1 Baseline Instruction Dataset
Table 11 displays the data sources that are used to prepare the dataset described in Section 3.2. These data are pooled and preprocessed into instruction-program pairs by stripping away Markdown formatting and natural language explanations from completions (Figure 9 and 10). In our experiments, we use the resultant data to finetune baseline models, comparing their performance to those of LMs fine-tuned on edit sequences generated with LintSeq from the same set of instruction-program pairs.
@@ E.1 Baseline Instruction Dataset
E.2 Procedures and Hyperparameters
We instruction finetune all models with Microsoft DeepSpeed using the ZeRO++ protocol for stage three sharding. For the largest of these models, we also use CPU parameter offloading to accelerate experiments (Wang et al., 2023a; Ren et al., 2021). When fine-tuning models on LintSeq data, we add a new token “<|diff|>” to tokenizers (Section 2.5) and resize model embeddings accordingly.
@@ E.2 Procedures and Hyperparameters
Across all of the fine-tuning experiments conducted in this paper, we train model-data variants with the same batch size and for an equal number of total optimizer steps. This optimizer step count corresponds to ten epochs of fine-tuning with the baseline instruction tuning dataset described in Section 3.2. We save intermediate checkpoints at equal optimizer step intervals in all experiments, and we report benchmark scores for the best performing checkpoint from each model-data variant.
@@
Table 13: All other instruction fine-tuning settings, re-used across experiments.
Appendix F More on Synthetic Data Generation with LintSeq
F.1 Examples of Generated Synthetic Edit Trajectories
@@
F.2 Tuning LintSeq Example Count